# Holographic inflation and holographic dark energy from entropy of the anti-de Sitter black hole

### v1

@@ Abstract
Based on the entropy of anti-de Sitter black hole, a new holographic dark energy model has been proposed. When the Hubble horizon and particle horizon are chosen as the IR cutoff, the late-time accelerated expansion of universe is realized. In this paper, we consider the Hubble horizon as the IR cutoff to investigate holographic inflation and slow-roll inflation in this model. We find that slow-roll inflation with the chaotic potential $V_{0}\phi^{n}$ is favored by Planck results for some special cases, such as $n=1/3$ and $n=1/2$, while holographic inflation is not supported by Planck results. Then, we use the dynamical analysis method, statefinder diagnostic pairs, and the Hubble diagram to analyze this model. Our results indicate that when $b^{2}$ takes a small value, this model cannot be distinguished from the standard $\Lambda$CDM model and can serve as an alternative to it.
@@ I Introduction
This paper has three main goals: to investigate inflation, to address the unresolved question regarding classical instability, and to examine the evolution of the universe in NHDEH. The paper is organized as follows: In Section II, we will examine inflation in NHDEH. In Section III, we will analyze both the evolution of the universe and the classical instability in NHDEH. Finally, our main conclusions are presented in Section IV.
@@ II Inflation
Here, $\kappa^{2}=8\pi G$, $b$ is a dimensionless parameter, and $\Lambda$ represents the cosmological constant, which provides a possibility of obtaining the late-time acceleration of the universe. For $\Lambda=0$, the energy density of NHDEH will be reduced to that of the original holographic dark energy Hsu2004 .
@@ II.2 Slow-roll inflation
Although slow-roll inflation can be realized in NHDEH, it is uncertain whether we can achieve the required e-folding number $N$, i.e., $7050<N<$ 70. To answer this question, we will conduct a further analysis of slow-roll inflation in NHDEH. According to the definition of the e-folding number $N$, we can write it as
@@
Here, Eq. (18) is used, the subscript e denotes when inflation ended, while the subscript i represents when inflation began. Substituting Eqs. (18) and  (28) into Eq. (32), we get
@@
Since the first slow-roll parameter at the end of inflation is known to take the value $\epsilon_{e}\simeq 1$, we can solve the expression for $\phi_{e}$ using this relation. Then, by substituting $\phi_{e}$ into Eq. (32), the expression for $\phi_{i}$ can be obtained. Combining Eqs.  (29),  (30), and $\phi_{i}$, we can obtain the expressions for $n_{s}$ and $r$. Similar to the situation in holographic inflation, the expressions for $\phi_{e}$, $\phi_{i}$, $n_{s}$, and $r$ depend on the values of $n$, $b^{2}$, and $N$ and cannot be expressed analytically. We will use figures to show their relationships.
Fixing the values of $n$ and $b^{2}$, we plot the predictions of the chaotic potential  (28) in the $r_{0.002}-n_{s}$ plane in Fig. (3), where we overlay our analytical results with Planck 2018 data. These figures show that slow-roll inflation can be realized for $n=\frac{1}{3}$, $\frac{1}{2}$, and $\frac{2}{3}$ by selecting the appropriate values for $b^{2}$ and the e-folding number $N$. The value of $r_{0.002}$ decreases as $N$ increases from $50$ to $70$ and as $b^{2}$ increases, while $n_{s}$ decreases as $N$ decreases from $70$ to $50$ and as $b^{2}$ decreases. With this comparison to the Planck 2018 data, the results favor the cases where $n=\frac{1}{3}$ with $b^{2}=0.50$, and $n=\frac{1}{2}$ with $b^{2}=0.65$. It is clear that an increased value of $n$ may result in a mismatch with the observations. In addition, using the binned Pantheon, cosmic chronometers, and baryon acoustic oscillations datasets, the observed value for $b^{2}$ is $b^{2}=0.01$ Nakarachinda2022 . This value is not excluded when $n=\frac{1}{3}$. Notably, the chaotic potential (28) is ruled out by the Planck 2018 data in the context of standard slow-roll inflation Planck2020a . However, slow-roll inflation is realized within NHDEH and is consistent with the Planck 2018 data.
@@ III Evolution of universe
In the previous section, we have analyzed holographic inflation and slow-roll inflation in NHDEH. After inflation ends, i.e., $\epsilon_{e}\simeq 1$, the universe reheats, after which the evolution of the hot ”big bang” universe begins. During the reheating phase, the inflaton begins to oscillate around the bottom of the potential and is expected to decay into other particles Bassett2006 ; Cai2015 . Then, the universe evolves from a radiation-dominated epoch, transitions into a matter-dominated epoch, and eventually evolves into an epoch dominated by dark energy.To discuss the evolution of the universe after inflation ends, we assume the scalar field $\phi$ decays into radiation and pressureless matter. Thus, the first Friedmann equation  (3) becomes
@@
III.1 Evolution and stability
To analyze the evolution of the universe in NHDEH, we take $\Omega^{0}_{r}=0.0001$, $\Omega^{0}_{m}=0.3111$, $\Omega^{0}_{D}=0.6888$, and $H_{0}=67.66kms^{-1}Mpc^{-1}$ Planck2020 as the initial conditions throughout this paper. Then, solving Eqs. (40) and (41) numerically, we obtain the evolution curves of $\Omega_{de}$, $\Omega_{m}$, $\omega_{de}$, and $q$, which are plotted in Figs. (4) and  (5). From Fig. 4, we can see that $\Omega_{de}$ approaches $1$ and $\Omega_{m}$ approaches 00 as the redshift decreases, and the universe can be dominated by NHDEH during its late-time evolution. However, in the relatively early stages of evolution, as $b^{2}$ increases, $\Omega_{de}$ increases while $\Omega_{m}$ decreases, indicating that a smaller $b^{2}$ is required to accurately describe the evolution of the universe in NHDEH. These results indicate that a small $b^{2}$ can be utilized to describe the evolution of the universe. The left panel of Fig. 5 shows that the NHDEH behaves as quintessence, and $\omega_{de}$ approaches $-1$, behaving like a cosmological constant during the late-time evolution. The right panel of Fig. 5 indicates that the late-time acceleration can be achieved, and a suitable range for the transition redshift($0.48\leq z_{t}<1$) is obtainable. In addition, the left panel of Fig. 6 shows that this model is stable and the squared sound speed $v^{2}_{s}$ is independent of $b^{2}$.

### v2

@@ Abstract
Based on the entropy of anti-de Sitter black hole, a new holographic dark energy model has been proposed. When the Hubble horizon and particle horizon are chosen as the IR cutoff, the late-time accelerated expansion of universe is realized. In this paper, we consider the Hubble horizon as the IR cutoff to investigate holographic inflation and slow-roll inflation in this model. We find that slow-roll inflation with the chaotic potential $V_{0}\phi^{n}$ is favored by Planck results for some special cases, such as $n=1/3$ and $n=1/2$, while holographic inflation is not supported by Planck results. Then, we analyze the reheating temperature and the number of reheating e-folds in this model, and we find that the results favor the cases $n=1/3$ and $n=1/2$. Finally, we use the dynamical analysis method, statefinder diagnostic pairs, and the Hubble diagram to analyze this model. Our results indicate that when $b^{2}$ takes a small value, this model cannot be distinguished from the standard $\Lambda$CDM model and can serve as an alternative to it.
@@ I Introduction
After the end of inflation, the universe enters a reheating phase, during which the inflaton field oscillates around the minimum of its potential and gradually decays into other particles Bassett2006 ; Cai2015 . In this phase, radiation and matter are produced through the decay of the inflaton field or other fields. These particles interact and eventually thermalize, reaching equilibrium at a reheating temperature $T_{reh}$. During this phase, the universe must thermalize at an extremely high temperature to facilitate big bang nucleosynthesis and baryogenesis. As a result, the temperature of the universe rises, leading to the radiation-dominated era and ultimately transitioning into the standard hot big bang model. To characterize the reheating phase, several important reheating parameters have been proposed: the reheating temperature $T_{reh}$, which is constrained by big bang nucleosynthesis Steigman2007 and the energy scale of inflation, typically lies in the range $10^{-2}GeV\leq T_{reh}\leq 10^{16}GeV$; the effective equation of state parameter $\omega_{reh}$, often assumed to be constant, is bounded within the range $-\frac{1}{3}\leq\omega_{reh}\leq 1$ Dai2014 ; Cook2015 ; Munoz2015 ; Goswami2018 ; Zhou2022 ; Yadav2024 ; Zhang2021 ; Zhang2023 ; Martin2015 ; the number of reheating e-folds $N_{reh}$, defined as the number of e-folds from the end of inflation to the beginning of the radiation-dominated era, is influenced by the effective equation of state parameter $\omega_{reh}$ and the inflationary model Dai2014 ; Cook2015 ; Munoz2015 ; Goswami2018 ; Zhou2022 ; Yadav2024 ; Zhang2021 ; Zhang2023 . To describe the reheating phase in the early universe, several reheating models have been proposed, including the perturbative decay of an oscillating inflaton field at the end of inflation Dolgov1982 ; Abbott1982 , as well as non-perturbative processes such as parametric resonance decay or tachyonic instability Traschen1990 ; Shtanov1995 ; Bassett2006 ; Fu2017 ; Fu2019a ; Jin2020 ; Li2020 .
@@
This paper has four main objectives: to investigate inflation, to explore reheating, to address the unresolved question of classical instability, and to examine the evolution of the universe in NHDEH. The paper is organized as follows: In Section II, we examine inflation in NHDEH. In Section III, we explore reheating in NHDEH. In Section IV, we analyze both the evolution of the universe and the classical instability in NHDEH. Finally, our main conclusions are presented in Section V.
@@ II Inflation
Here, $\kappa^{2}=8\pi G$, $b$ is a dimensionless parameter, and $\Lambda$ represents the cosmological constant, which provides a possibility of obtaining the late-time acceleration of the universe. For $\Lambda=0$, the energy density of NHDEH reduces to that of the original holographic dark energy Hsu2004 , which cannot explain the late-time cosmic acceleration or describe the complete evolutionary history of the universe.
@@ II.2 Slow-roll inflation
Although slow-roll inflation can be realized in NHDEH, it is uncertain whether we can achieve the required e-folds number $N$, i.e., $7050<N<$ 70. To answer this question, we will conduct a further analysis of slow-roll inflation in NHDEH. According to the definition of the e-folds number $N$, we can write it as
@@
Here, Eq. (18) is used, the subscript e denotes the time when inflation ended, while the subscript i represents the time when inflation began. Then, substituting Eqs. (18) and  (28) into Eq. (31), we get
@@
Since the first slow-roll parameter at the end of inflation is known to take the value $\epsilon_{e}\simeq 1$, we can solve the expression for $\phi_{e}$ using this relation. By substituting $\phi_{e}$ into Eq. (32), the expression for $\phi_{i}$ can be obtained. Then, combining Eqs. (29),  (30), and $\phi_{i}$, we can obtain the expressions for $n_{s}$ and $r$. Similar to the situation in holographic inflation, the expressions for $\phi_{e}$, $\phi_{i}$, $n_{s}$, and $r$ depend on the values of $n$, $b^{2}$, and $N$ and cannot be expressed analytically. We will use figures to show their relationships.
Fixing the values of $n$ and $b^{2}$, we plot the predictions of the chaotic potential  (28) in the $r_{0.002}-n_{s}$ plane in Fig. (3), where we overlay our analytical results with Planck 2018 data. These figures show that slow-roll inflation can be realized for $n=\frac{1}{3}$, $\frac{1}{2}$, and $\frac{2}{3}$ by selecting the appropriate values for $b^{2}$ and the e-folds number $N$. The value of $r_{0.002}$ decreases as $N$ increases from $50$ to $70$ and as $b^{2}$ increases, while $n_{s}$ decreases as $N$ decreases from $70$ to $50$ and as $b^{2}$ decreases. With this comparison to the Planck 2018 data, the results favor the cases where $n=\frac{1}{3}$ with $b^{2}=0.50$, and $n=\frac{1}{2}$ with $b^{2}=0.65$. It is clear that an increased value of $n$ may result in a mismatch with the observations. In addition, using the binned Pantheon, cosmic chronometers, and baryon acoustic oscillations datasets, the observed value for $b^{2}$ is $b^{2}=0.01$ Nakarachinda2022 . This value is not excluded when $n=\frac{1}{3}$. Notably, the chaotic potential (28) is ruled out by the Planck 2018 data in the context of standard slow-roll inflation Planck2020a . However, slow-roll inflation is realized within NHDEH and is consistent with the Planck 2018 data.
@@ III Reheating
In the previous section, we have analyzed holographic inflation and slow-roll inflation in NHDEH. After inflation ends, i.e., $\epsilon_{e}\simeq 1$, the universe reheats, after which the evolution of the hot big bang universe begins. During the reheating phase, the inflaton begins to oscillate around the bottom of the potential and is expected to decay into other particles Bassett2006 ; Cai2015 . To describe the reheating phase, the reheating temperature $T_{reh}$, the effective equation of state parameter $\omega_{reh}$ and the number of reheating e-folds $N_{reh}$ are introduced Dai2014 ; Cook2015 ; Munoz2015 ; Goswami2018 ; Zhou2022 ; Yadav2024 ; Zhang2021 ; Zhang2023 . In this section, we will discuss whether $T_{reh}$ and $N_{reh}$ produced during reheating in NHDEH can be supported by observations.
Combining Eqs. (17),  (18), and  (25), the energy density and Hubble parameter during inflation can be written as
$\rho_{\phi}=\frac{3V}{3-\epsilon}+\frac{3b^{2}\Lambda\epsilon}{\kappa^{2}(3-\epsilon)},$ (33)
$H^{2}=\frac{\kappa^{2}V+3b^{2}\Lambda}{(1-b^{2})(3-\epsilon)}.$ (34)
Here, the potential $V$ is given in Eq. (28). After inflation ends, the slow-roll parameter $\epsilon$ satisfies $\epsilon_{e}\simeq 1$. Hence, the energy density at the end of inflation becomes
$\rho_{e}=\frac{3}{2}\Big{(}V_{e}+\frac{b^{2}V}{\kappa^{2}}\Big{)}.$ (35)
Similar to the definition of the e-folds number during inflation, the number of e-folds between the time when a mode $k$ crosses the Hubble horizon and the end of inflation is defined as
[…]
where the subscript k denotes the time when the mode $k$ crosses the Hubble radius. And the number of reheating e-folds $N_{reh}$ is given by
$\Delta N_{reh}=\ln\Big{(}\frac{a_{reh}}{a_{e}}\Big{)},$ (37)
where the subscript reh denotes the end of reheating. Since $N_{reh}$ is defined as the number of e-folds from the end of inflation to the beginning of the radiation-dominated era, it encodes both the epoch of reheating and the subsequent thermalization process. The subsequent evolution of the universe is governed by its energy density
$\rho_{reh}=\frac{\pi^{2}}{30}g_{reh}T^{4}_{reh},$ (38)
where $g_{reh}$ represents the effective number of relativistic species and $T_{reh}$ is the temperature at the end of reheating. Then, considering the continuity equation, we have
$\dot{\rho}+3H(\rho+p)=\dot{\rho}+3H(1+\omega_{reh})\rho=0.$ (39)
In view of this equation, we get Goswami2018
$\rho_{reh}=\rho_{e}e^{-3N_{reh}(1+\overline{\omega}_{reh})}.$ (40)
Here, $\overline{\omega}_{reh}$ is the average equation of state parameter during reheating Martin2015 . Eq. (40) can be rewritten as
$\frac{a_{reh}}{a_{e}}=e^{N_{reh}}=\Big{(}\frac{\rho_{reh}}{\rho_{e}}\Big{)}^{-\frac{1}{3(1+\overline{\omega}_{reh})}}.$ (41)
In order to relate the observed wavenumber of any physical scale today $\frac{k}{a_{0}}$ to its value at the time of Hubble crossing during inflation $\frac{k}{a_{k}}$, we have Goswami2018
$\frac{k}{a_{k}}=\frac{k}{a_{0}}\frac{a_{0}}{a_{k}}=\frac{k}{a_{0}}\frac{a_{0}}{a_{eq}}\frac{a_{eq}}{a_{reh}}\frac{a_{reh}}{a_{e}}\frac{a_{e}}{a_{k}}=\frac{k}{a_{0}}(1+z_{eq})\Big{(}\frac{\rho_{reh}}{\rho_{eq}}\Big{)}^{\frac{1}{4}}e^{N_{reh}}e^{\Delta N_{k}},$ (42)
where the subscript eq denotes the epoch of matter-radiation equality. Combining Eqs. (35),  (41), and  (42), we obtain
$\frac{k}{a_{k}}=H_{k}=\frac{k}{a_{0}}(1+z_{eq})\rho_{reh}^{\frac{3\overline{\omega}_{reh}-1}{12(1+\overline{\omega}_{reh})}}\rho_{eq}^{-\frac{1}{4}}\Big{[}\frac{3}{2}\Big{(}V_{e}+\frac{b^{2}\Lambda}{\kappa^{2}}\Big{)}\Big{]}^{\frac{1}{3(1+\overline{\omega}_{reh})}}e^{\Delta N_{k}}.$ (43)
Substituting Eq. (38) into  (43), the expression for $T_{reh}$ can be written as
$\ln(T_{reh})$ $=\frac{3(1+\overline{\omega}_{reh})}{3\overline{\omega}_{reh}-1}\Big{[}\ln(H_{k})-\ln\Big{(}\frac{k}{a_{0}}\Big{)}-\ln(1+z_{eq})-\Delta N_{k}+\ln\Big{(}\rho_{eq}^{\frac{1}{4}}\Big{)}\Big{]}$ (44)
$-\frac{1}{3\overline{\omega}_{reh}-1}\ln\Big{[}\frac{3}{2}\Big{(}V_{e}+\frac{b^{2}\Lambda}{\kappa^{2}}\Big{)}\Big{]}-\frac{1}{4}\ln\Big{(}\frac{\pi^{2}}{30}g_{reh}\Big{)}.$
Combining Eqs. (38),  (41), and  (44), we can obtain the expression for $N_{reh}$ as follows
$N_{reh}$ $=\frac{4}{3\overline{\omega}_{reh}-1}\Big{[}\ln\Big{(}\frac{k}{a_{0}}\Big{)}+\ln(1+z_{eq})+\Delta N_{k}-\ln(H_{k})-\ln\Big{(}\rho_{eq}^{\frac{1}{4}}\Big{)}\Big{]}$ (45)
$+\frac{1}{3\overline{\omega}_{reh}-1}\ln\Big{[}\frac{3}{2}\Big{(}V_{e}+\frac{b^{2}\Lambda}{\kappa^{2}}\Big{)}\Big{]}.$
Eqs. (44) and  (45) are the two key relationships that can be used to parameterize reheating in slow-roll inflationary models. To study the constraints on reheating from the CMB data, we consider a pivot scale $k_{*}$ which corresponds to the time when the observable CMB scales exit the Hubble radius $H_{*}$ during inflation. For the pivot scale $k_{*}$, we take the value $\frac{k_{*}}{a_{0}}=0.05Mpc^{-1}$ used by the Planck collaboration. For $k=k_{*}$, the amplitude of scalar power spectra is given in terms of $H_{*}$ as
$A_{s}=P_{\zeta}(k_{*})=\frac{\kappa^{2}H^{2}_{*}}{8\pi^{2}\epsilon_{*}},$ (46)
which gives the relation
$H_{*}=\frac{\pi}{\kappa}\sqrt{8A_{s}\epsilon_{*}}.$ (47)
The number of e-folds remaining after the pivot scale $k_{*}$ crosses the Hubble radius is given by
[…]
where $\phi_{e}$ can be solved numerically using Eqs. (17),  (18), and  (25) under the condition $\epsilon_{e}\simeq 1$. Then, using Eqs. (17),  (18), and  (25), the spectral index $n_{s}$ (29) can be rewritten as
$n_{s}=1-\Big{[}2\Big{(}1+2b^{2}+\frac{2}{n}\Big{)}-12\frac{b^{2}(n-1)}{\kappa^{2}n}\Big{]}\epsilon_{*}.$ (49)
Using Eq. (47), we obtain the relation
$V_{e}+\frac{b^{2}\Lambda}{\kappa^{2}}=\frac{3H^{2}_{*}\phi^{n}_{e}}{\kappa^{2}\phi^{n}_{*}}+\frac{b^{2}}{\kappa^{2}}\Big{[}\Lambda-3(\Lambda+H^{2}_{*})\frac{\phi^{n}_{e}}{\phi^{n}_{*}}\Big{]}.$ (50)
Now, using Eqs. (47),  (48),  (49), and  (50), the reheating temperature $T_{reh}$ (44) and the number of reheating e-folds $N_{reh}$ (45) can be expressed as functions of $\overline{\omega}_{reh}$ and $n_{s}$ for the pivot scale. Then, taking $\hbar=c=1$, $g_{reh}=100$, $\kappa=4.106\times 10^{-9}GeV^{-1}$ and using Planck 2018 results of $A_{s}=2.105\times 10^{9}$, $z_{eq}=3387$, and $H_{0}=67.66kms^{-1}Mpc^{-1}$ Planck2020 , we have plotted the figures for $T_{reh}$ and $N_{reh}$ versus $n_{s}$, which are shown in Figs. (4) and  (5), respectively. Here, we adopt $b^{2}=0.01$ Nakarachinda2022 . The blue and light blue shaded regions correspond to the $1\sigma$ and $2\sigma$ bounds on $n_{s}$ from Planck 2018 data (TT,TE,EE+lowE+lensing+BK15+BAO) Planck2020 .
From Figs. (4) and  (5), it can be observed that the instantaneous reheating is excluded by Planck results in NHDEH, as the corresponding spectral index $n_{s}$ exceeds the $2\sigma$ bound on $n_{s}$ from Planck data. In Fig. (4), within the $2\sigma$ bound on $n_{s}$ from Planck results, the data favor a negative $\overline{\omega}_{reh}$, and the possible allowed values of reheating temperature $T_{reh}$ are greater than $10^{-2}GeV$. Curves for $\overline{\omega}_{reh}=-\frac{1}{3}$ predict the allowed values of reheating temperature $T_{reh}$ to be $4.43\times 10^{-1}GeV$ to $1.17\times 10^{5}GeV$ for $n=\frac{1}{3}$ and $1.72\times 10^{-2}GeV$ to $5.68\times 10^{2}GeV$ for $n=\frac{1}{2}$. Fig. (5) shows that the number of reheating e-folds $N_{reh}$ are greater than $9$ for $n=\frac{1}{3}$ and greater than $20$ for $n=\frac{1}{2}$ within the $2\sigma$ bound on $n_{s}$ from Planck results. Curves for $\overline{\omega}_{reh}=-\frac{1}{3}$ predict $359<N_{reh}<$ 35 for $n=\frac{1}{3}$ and $4220<N_{reh}<$ 42 for $n=\frac{1}{2}$. Thus, under the constraint $b^{2}=0.01$ Nakarachinda2022 , the results from reheating favor the cases $n=\frac{1}{3}$ and $n=\frac{1}{2}$.
Figure 4: Reheating temperature $T_{reh}$ versus $n_{s}$ for NHDEH. The left panel is plotted for $n=\frac{1}{3}$, while the right one is for $n=\frac{1}{2}$.
Figure 5: Number of reheating e-folds $N_{reh}$ versus $n_{s}$ for NHDEH. The left panel is plotted for $n=\frac{1}{3}$, while the right one is for $n=\frac{1}{2}$.
For the case $b=0$ and $n=2$, NHDEH reduces to the quadratic chaotic inflationary model $V=\frac{1}{2}m^{2}\phi^{2}$, and our results for $T_{reh}$ and $N_{reh}$ will reduce to those in  Dai2014 ; Goswami2018 ; Yadav2024 .
IV Evolution of universe
In the previous section, we discussed reheating in NHDEH. After reheating ends, the universe evolves from a radiation-dominated epoch, transitions to a matter-dominated epoch, and eventually enters an epoch dominated by dark energy. To discuss the evolution of the universe after inflation, we assume that the scalar field $\phi$ decays into radiation and pressureless matter. Thus, the first Friedmann equation  (3) becomes
@@ IV Evolution of universe
IV.1 Evolution and stability
To analyze the evolution of the universe in NHDEH, we take $\Omega^{0}_{r}=0.0001$, $\Omega^{0}_{m}=0.3111$, $\Omega^{0}_{D}=0.6888$, and $H_{0}=67.66kms^{-1}Mpc^{-1}$ Planck2020 as the initial conditions throughout this paper. Then, solving Eqs. (58) and (59) numerically, we obtain the evolution curves of $\Omega_{de}$, $\Omega_{m}$, $\omega_{de}$, and $q$, which are plotted in Figs. (6) and  (7). From Fig. (6), we can see that $\Omega_{de}$ approaches $1$ and $\Omega_{m}$ approaches 00 as the redshift decreases, and the universe can be dominated by NHDEH during its late-time evolution. However, in the relatively early stages of evolution, as $b^{2}$ increases, $\Omega_{de}$ increases while $\Omega_{m}$ decreases, indicating that a smaller $b^{2}$ is required to accurately describe the evolution of the universe in NHDEH. These results indicate that a small $b^{2}$ can be utilized to describe the evolution of the universe. The left panel of Fig. (7) shows that the NHDEH behaves as quintessence, and $\omega_{de}$ approaches $-1$, behaving like a cosmological constant during the late-time evolution. The right panel of Fig. (7) indicates that the late-time acceleration can be achieved, and a suitable range for the transition redshift($0.48\leq z_{t}<1$) is obtainable. In addition, the left panel of Fig. (8) shows that this model is stable and the squared sound speed $v^{2}_{s}$ is independent of $b^{2}$.